Following information is available in respect of a company named, as Beta Ltd
Particulars Amount
Profit before taxes Rs 50 lakhs
Dividend declared per share Rs2
Price of the share prevalent on stock exchange Rs 200
Applicable tax rate 35%
Share capital of company 5 lakh shares of Rs 10 each
Calculate-
• Earnings per share
• Dividend Pay-out Ratio
• Price earnings ratio
1
Expert's answer
2019-12-03T10:05:45-0500

Earnings per share = (Net Income- Preferred Dividends)/Weighted Average Shares Outstanding

EPS=(50*(1-0.35)*100000-2*5*100000)/(5*100000)=Rs 4.5

Dividend Pay-out Ratio = Dividends/Net income

DPR==2*5*100000/(50*(1-0.35)*100000)=0.307

Price earnings ratio= Price per share/Earnings per share

PER=200/4.5=44.44